OpenStudy (anonymous):

~~~Giving out a medal and fanning whoever helps me!!!~~~ Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

@mayaal

OpenStudy (anonymous):

@SA5UK3

OpenStudy (mayaal):

do you want the answer?

OpenStudy (anonymous):

@Nikki-Albarn09

OpenStudy (anonymous):

Yeah please i really dont know how to do this

OpenStudy (anonymous):

Do you know what a polynomial is? I would suggest doing a difference of squares and then having the bottom as one of the roots. When it simplifies it'll just be a simple linear function.

OpenStudy (mayaal):

2xsqrd+4x-6/2 =xsqrd+2x-3

OpenStudy (anonymous):

I dont really know what a polynomial is and im really confused of what you just said

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

\[\frac{ x^2-49 }{ (x+7) }\] Now do you know what the top breaks down into?

OpenStudy (anonymous):

here check this link out it will help http://answers.yahoo.com/question/index?qid=20100918071111AAqngOv

OpenStudy (mayaal):

a polynomial is an expression with different terms with different exponents

OpenStudy (anonymous):

Ok @mayaal but Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

Ohh ok

OpenStudy (anonymous):

Ok @Nikki-Albarn09

OpenStudy (anonymous):

@love_jessika15

OpenStudy (mayaal):

divide the quotient by each term in the expression,by 2xsqrd,4x,and 6.

OpenStudy (anonymous):

uh huh then whats another way?

OpenStudy (mayaal):

the other way is by multiplying the simplified expression by the quotient

OpenStudy (mayaal):

multiply x squared+2x-3 by 2,and you will get 2x squared+4x-6

OpenStudy (mayaal):

do u get it?

OpenStudy (mayaal):

?????????????

OpenStudy (anonymous):

Lol yeah i get it but how you would check your quotient for accuracy.

OpenStudy (mayaal):

i told u.multiply the divisor by the quotient

OpenStudy (mayaal):

multiply x squared+2x-3 by 2,and you will get 2x squared+4x-6

OpenStudy (anonymous):

Ohhh

OpenStudy (mayaal):

r u good?

OpenStudy (anonymous):

Uhh i think give me one second

OpenStudy (anonymous):

2x^2+4x-6/2 =x^2+2x-3 divide the quotient by each term in the expression,by 2x^2,4x,and 6.The other way is by multiplying the simplified expression by the quotient.Multiply x^2+2x-3 by 2,and you will get 2x^2+4x-6.Multiply the divisor by the quotient in order to check the accuracy.

OpenStudy (anonymous):

Is that a good explanation for my question?

OpenStudy (mayaal):

multiplying the simplified expression by the divisor

OpenStudy (mayaal):

not the quotient

OpenStudy (anonymous):

Ok so change the last sentence to that?

OpenStudy (anonymous):

@mayaal

OpenStudy (mayaal):

no,the 2nd sentence